# RAEmoLLM: Retrieval Augmented LLMs for Cross-Domain Misinformation Detection Using In-Context Learning based on Emotional Information

### v1

@@ Abstract
Misinformation is prevalent in various fields such as education, politics, health, etc., causing significant harm to society. However, current methods for cross-domain misinformation detection rely on time and resources consuming fine-tuning and complex model structures. With the outstanding performance of LLMs, many studies have employed them for misinformation detection. Unfortunately, they focus on in-domain tasks and do not incorporate significant sentiment and emotion features (which we jointly call affect). In this paper, we propose RAEmoLLM, the first retrieval augmented (RAG) LLMs framework to address cross-domain misinformation detection using in-context learning based on affective information. It accomplishes this by applying an emotion-aware LLM to construct a retrieval database of affective embeddings. This database is used by our retrieval module to obtain source-domain samples, which are subsequently used for the inference module’s in-context few-shot learning to detect target domain misinformation. We evaluate our framework on three misinformation benchmarks. Results show that RAEmoLLM achieves significant improvements compared to the zero-shot method on three datasets, with the highest increases of 20.69%, 23.94%, and 39.11% respectively. This work will be released on https://github.com/lzw108/RAEmoLLM.
@@
Zhiwei Liu1thanks: Work done during Zhiwei’s internship at the University of Melbourne Kailai Yang1 Qianqian Xie2  Christine de Kock3 Sophia Ananiadou1,4,5Eduard Hovy3,6 1 The University of Manchester  2 The Fin AI  3 The University of Melbourne 4 Artificial Intelligence Research Center  5 Archimedes/Athena RC  6 Carnegie Mellon University {zhiwei.liu,kailai.yang,sophia.ananiadou}@manchester.ac.uk xqq.sincere@gmail.com,{christine.dekock,eduard.hovy}@unimelb.edu.au
@@ 1 Introduction
Although some studies address cross-domain misinformation detection Comito et al. (2023); Tang et al. (2023); Shi et al. (2023), they require time-consuming fine-tuning, and apply only traditional machine learning methods or complex deep learning methods. Recently, LLMs have achieved impressive results in various tasks through zero-shot, few-shot Li (2023), or instruction tuning Zhang et al. (2023a). Many researchers have applied LLMs to identify misinformation Li et al. (2023); Hu et al. (2024); Cheung and Lam (2023). However, these methods perform only in-domain misinformation detection. Moreover, emotions and sentiments (which we jointly call affect) are important characteristics of human expression and communication Hakak et al. (2017). When authors publish misinformation, they often consciously choose specific emotions to capture the attention and resonance of readers to encourage rapid spread Keen (2006); Liu et al. (2024d). Unfortunately, there are few LLMs that utilize affective information to detect misinformation, and the only ConspEmoLLM Liu et al. (2024b) are developed based on an emotional LLM, which does not make full use of affective information, has no cross-domain ability, and also needs time-consuming fine-tuning.
@@
To address these issues, we propose the first retrieval augmented (RAG) LLMs framework based on emotional information (RAEmoLLM), to address cross-domain misinformation detection using in-context learning based on affective information. RAEmoLLM contains three modules: (1) In the index construction module, we apply EmoLLaMA-chat-7B Liu et al. (2024c) to encode all domain corpora, obtaining implicit embeddings to construct the retrieval database as well as explicit affective labels. We also conduct a comprehensive affective analysis to demonstrate the effectiveness of affective information for discriminating between true and misinformation. (2) The retrieval module recommends the top K affect-related examples (text-label pairs) from the source domain corpus according to the target domain content, obtained from the retrieval database. (3) These examples are utilized as the few-shot demonstrations in the inference module, which is driven by a prompt template to guide the LLM to verify the target content for misinformation. The template helps combine implicit and explicit affective information.
@@
  • We propose RAEmoLLM, the first framework for cross-domain misinformation detection using ICL based on affective information, which does not require fine-tuning. Experimental results show that RAEmoLLM outperforms the zero-shot methods and the few-shot methods without using affective information.
  • We evaluate RAEmoLLM on a variety of misinformation benchmarks, including fake news, rumors, and conspiracy theory datasets. Results show that RAEmoLLM achieves significant improvements compared to the zero-shot method on three datasets, with the highest increases of 20.69%, 23.94%, and 39.11% respectively, which illustrate the effectiveness of RAEmoLLM framework.
2 Related Work
2.1 Misinformation detection
Cross-domain misinformation detection: Comito et al. (2023) propose a deep learning-based architecture able to mitigate this problem by yielding high-level cross-domain features. Tang et al. (2023) design the News Optimal Transport to learn transferable features across domains by aligning the source and target news using Optimal Transport (OT) techniques. Shi et al. (2023) develop a rough-fuzzy graph learning framework that uses representations of cross-domain sample uncertainty structural information, and captures shared general features across domains. But these methods all require complex structures and fine-tuning strategies.
Based on affective information: Emotion and sentiment are important features for misinformation detection Liu et al. (2024d). Zhang et al. (2023b) combine the use of semantic and sentiment information, along with propagation information for rumor detection. Dong et al. (2022a) design a sentiment-aware hyper-graph attention network for fake news detection. Liu et al. (2024b) develop a conspiracy theory detection LLM by fine-tuning EmoLLaMA Liu et al. (2024c). Choudhry et al. (2022) utilize emotional information for fake news detection based on an adversarial learning structure. Unfortunately, these works either have complex structural designs or fine-tuned models, which require significant time and computational resources. The RAEmoLLM in this article applies the ICL method based on affective information, which has a simple structure and does not involve fine-tuning.
Figure 1: The architecture of RAEmoLLM. D: Domain. T: Target domain. S: Source domain. C: Corpus. L: Label. Aff: Affective information.
2.2 In-context learning
Liu et al. (2024a) develop in-context curriculum learning, a simple but helpful demonstration ordering method for ICL that gradually increases the complexity of prompt demonstrations. Xu and Zhang (2024) propose in-context reflection to strategically select demonstrations that reduce the discrepancy between the LLM’s outputs and the actual input-output mappings. Long et al. (2023) propose a retrieval-enhanced language model to address cross-domain problems, in which they train language models by learning both target domain distribution and the discriminative task signal simultaneously with the augmented cross-domain in-context examples. Inspired by this work, we propose the RAEmoLLM.
@@ 3 Methodology
This section introduces our method of cross-domain misinformation detection, using the index construction module, retrieval module and inference module. The overall architecture of RAEmoLLM is shown in Figure 1. In the index construction module we collect domain datasets, and employ an emotional LLM to obtain embeddings as well as affective labels to conduct a comprehensive affective analysis on them to detect the affective differences between real and false information. The implicit embeddings are adopted to construct the retrieval database. This database is used by the retrieval module to obtain source-domain examples. These results are the few-shot examples used by the inference module’s in-context learning to detect target domain misinformation.
@@ 3.1.1 Datasets
Table 2: Statistics values of EIreg and Vreg on different datasets. The t-test is conducted between legit/non-rumours/related and fake/rumours/conspiracy.

### v2

@@ Abstract
Misinformation is prevalent in various fields such as education, politics, health, etc., causing significant harm to society. However, current methods for cross-domain misinformation detection rely on effort- and resource-intensive fine-tuning and complex model structures. With the outstanding performance of LLMs, many studies have employed them for misinformation detection. Unfortunately, they focus on in-domain tasks and do not incorporate significant sentiment and emotion features (which we jointly call affect). In this paper, we propose RAEmoLLM, the first retrieval augmented (RAG) LLMs framework to address cross-domain misinformation detection using in-context learning based on affective information. RAEmoLLM includes three modules. (1) In the index construction module, we apply an emotional LLM to obtain affective embeddings from all domains to construct a retrieval database. (2) The retrieval module uses the database to recommend top K examples (text-label pairs) from source domain data for target domain contents. (3) These examples are adopted as few-shot demonstrations for the inference module to process the target domain content. The RAEmoLLM can effectively enhance the general performance of LLMs in cross-domain misinformation detection tasks through affect-based retrieval, without fine-tuning. We evaluate our framework on three misinformation benchmarks. Results show that RAEmoLLM achieves significant improvements compared to the other few-shot methods on three datasets, with the highest increases of 15.64%, 31.18%, and 15.73% respectively. This project is available at https://github.com/lzw108/RAEmoLLM.
@@
Zhiwei Liu1  Kailai Yang1  Qianqian Xie2  Christine de Kock3 Sophia Ananiadou1thanks: Corresponding AuthorEduard Hovy3,4 1 The University of Manchester  2 School of Artificial Intelligence, Wuhan University 3 The University of Melbourne  4 Carnegie Mellon University {zhiwei.liu,kailai.yang,sophia.ananiadou}@manchester.ac.uk xqq.sincere@gmail.com,{christine.dekock,eduard.hovy}@unimelb.edu.au
@@ 1 Introduction
Although some studies address cross-domain misinformation detection Comito et al. (2023); Tang et al. (2023); Shi et al. (2023), they require effort-intensive fine-tuning, and apply only traditional machine learning methods or complex deep learning methods. Recently, LLMs have achieved impressive results in various tasks through zero-shot, few-shot Li (2023), or instruction tuning Zhang et al. (2023a). Many researchers have applied LLMs to identify misinformation Li et al. (2023); Hu et al. (2024); Cheung and Lam (2023). However, these methods perform only in-domain misinformation detection. Moreover, emotions and sentiments (which we jointly call affect) are important characteristics of human expression and communication Hakak et al. (2017). When authors publish misinformation, they often consciously choose specific emotions to capture the attention and resonance of readers to encourage rapid spread Keen (2006); Liu et al. (2024d). Unfortunately, there are few LLMs that utilize affective information to detect misinformation, and the only ConspEmoLLM Liu et al. (2024b) are developed based on an emotional LLM, which does not make full use of affective information, has no cross-domain ability, and also needs time-consuming fine-tuning.
@@
To address these issues, we propose the first retrieval augmented (RAG) LLMs framework based on emotional information (RAEmoLLM), to address cross-domain misinformation detection using in-context learning based on affective information. RAEmoLLM contains three modules: (1) In the index construction module, we apply EmoLLaMA-chat-7B Liu et al. (2024c) to encode all domain corpora, obtaining implicit affective embeddings to construct the retrieval database as well as explicit affective labels. We also conduct a comprehensive affective analysis to demonstrate the effectiveness of affective information for discriminating between true and misinformation. (2) The retrieval module recommends the top K affect-related examples (text-label pairs) from the source domain corpus according to the target domain content, obtained from the retrieval database. (3) These examples are utilized as the few-shot demonstrations in the inference module, which is driven by a prompt template to guide the LLM to verify the target content for misinformation. The template helps combine implicit and explicit affective information. This framework effectively enhances the capabilities of LLMs in multiple cross-domain misinformation detection tasks through leveraging affective information, without the need for fine-tuning.
@@
  • We propose RAEmoLLM, the first framework for cross-domain misinformation detection using ICL based on affective information, which does not require fine-tuning. Experimental results show that RAEmoLLM outperforms the zero-shot method and other few-shot methods.
  • We evaluate RAEmoLLM on a variety of misinformation benchmarks, including fake news, rumours, and conspiracy theory datasets. Results show that RAEmoLLM achieves significant improvements compared to the other few-shot methods on three datasets, with the highest increases of 15.64%, 31.18%, and 15.73% respectively, which illustrate the effectiveness of RAEmoLLM framework.
@@ 2 Methodology
Figure 1: The architecture of RAEmoLLM. D: Domain. T: Target domain. S: Source domain. C: Corpus. L: Label. Aff: Affective information. M: Number of source domain data. Index Construction Module: Apply an emotional LLM to obtain affective embeddings to construct a retrieval database. Retrieval Module: Recommend top K examples (text-label pairs) from source domain data. Inference Module: Adopt the recommended examples as demonstrations for inference.
This section introduces our method of cross-domain misinformation detection, using the index construction module, retrieval module and inference module. The overall architecture of RAEmoLLM is shown in Figure 1. In the index construction module (Sec. 2.1), we collect domain datasets, and employ an emotional LLM to obtain affective embeddings as well as affective labels to conduct a comprehensive affective analysis on them to detect the affective differences between real and false information. The implicit embeddings are adopted to construct the retrieval database, which will be used by the retrieval module (Sec. 2.2) to obtain source-domain examples. These results are used as the few-shot examples for inference module’s (Sec. 2.3) in-context learning to detect target domain misinformation.
@@ 2.1 Index Construction Module
In this section, we first introduce the original datasets and the processing procedure at Sec. 2.1.1. We subsequently conduct affective analysis on these datasets and present how and why to obtain implicit and explicit affective information at Sec. 2.1.2. Finally, we apply the implicit affective information to construct the retrieval database (Sec. 2.1.3).
@@ 2.1.1 Datasets
Table 1: Statistics of datasets. AMTCele includes 7 domains. PHEME contains 9 domains (events). COCO has 12 domains (topics). For AMTCele and PHEME, we apply leave-one-domain-out strategy for evaluation. For COCO, we select 3 domains as test set.